OpenStudy (mathgenus123):

need help plz Factor the following expression 10x + 60

OpenStudy (mathgenus123):

dont answer just ex-plane how to solve it

563blackghost (563blackghost):

How many 10's can go into 60?

OpenStudy (mathgenus123):

6

OpenStudy (mathgenus123):

i dont know

OpenStudy (mathgenus123):

what is10x

563blackghost (563blackghost):

hmmm how many 10's go into 10?

OpenStudy (mathgenus123):

1

OpenStudy (mathgenus123):

i still am confused

563blackghost (563blackghost):

Its quite hard for me to explain this but I tried....the factorization would be.... \(\LARGE{10(x+6)}\) Because 10 is a factor of 60 and 10x...

563blackghost (563blackghost):

Another way is that you would divide by 10 to get x by itself to get the simplified equation.... \[\Large\frac{ 10x+60 }{ 10 }=x+6=10(x+6)\]

563blackghost (563blackghost):

Do you kinda get what I am saying?